the ratio of length of a school ground ti its width is 5:2 find the length if width is 40 meters
Answer by stanbon(75887)

the ratio of length of a school ground ti its width is 5:2 find the length if width is 40 meters
------
x/40 = 5/2
x = 200/2 = 100
length = 100 meters
====================
Cheers,
Stan H.
-----------------